OpenStudy (anonymous):

What is the 9th term of the geometric sequence where a1 = –5 and a6 = –5,120?

OpenStudy (anonymous):

–524,288 –458,752 –393,216 –327,680

OpenStudy (chaise):

The equation for any geometric sequence is: tn=ar^(n-1) -5=-5r(1-1) -5,120 = -5r^(-5120-1) Two equations with two unknown, you can solve them simultaneously Do you need help with this?